When the sum of 546 and three times a positive number is subtracted from the square of the number, the result is 154. Find the number.
Answer by nerdybill(7384)

When the sum of 546 and three times a positive number is subtracted from the square of the number, the result is 154. Find the number.
Let n = the number
then
n^2 - (546 + 3n) = 154
n^2 - 546 - 3n = 154
n^2 - 3n - 546 = 154
n^2 - 3n - 700 = 0
(n + 25)(n - 28) = 0
n = {-25, 28}
Since the problem specified a POSITIVE number:
n = 28